# Validation tests of Gaussian boson samplers with photon-number resolving detectors

### v1

@@ Abstract
An important challenge with the current generation of noisy, large-scale quantum computers is the question of validation. Does the hardware generate correct answers? If not, what are the errors? This issue is often combined with questions of computational advantage, but it is a fundamentally distinct issue. In current experiments, complete validation of the output statistics is generally not possible because it is exponentially hard to do so. Here, we apply phase-space simulation methods to partially verify recent experiments on Gaussian boson sampling (GBS) implementing photon-number resolving (PNR) detectors. The positive-P phase-space distribution is employed, as it uses probabilistic sampling to reduce complexity. It is $10^{18}$ times faster than direct classical simulation for experiments on $288$ modes where quantum computational advantage is claimed. When combined with binning and marginalization to improve statistics, multiple validation tests are efficiently computable, of which some tests can be carried out on experimental data. We show that the data as a whole shows discrepancies with theoretical predictions for perfect squeezing. However, a small modification of the GBS parameters greatly improves agreement. Hence, we suggest that such validation tests could form the basis of feedback methods to improve GBS quantum computer experiments.
@@ I Introduction
Theoretical and experimental interest in photonic networks employed as quantum computers has grown considerably [1, 2]. These devices interfere nonclassical photons in a large scale interferometric network, then measure the resulting output photons using photodetectors. The computational task is sampling from a computationally hard probability distribution. Gaussian boson sampling (GBS) has proved scalable to the sizes needed for quantum advantage, which recent experiments have now claimed [3, 4, 5, 6]. In this case, photons are prepared in a squeezed state, with output probabilities in a #P-hard computational class [7, 8]. For photon-number resolving (PNR) detectors this is a Hafnian, while for threshold detectors it is a Torontonian. Neither function is computable in polynomial time, so that the random counts cannot be classically replicated at large scale.
In this paper, we use positive P-representation simulations to validate Gaussian boson sampling with photon number resolving (PNR) detectors, extending the tests that can be applied to experimental data. This is essential, as a number of large scale GBS experiments have claimed quantum advantage using either PNR [5] or threshold detectors [3, 4, 6]. To perform the tests, we generalize a binning method developed by Mandel [9, 10] for simulating photon statistics of classical light. The original formulation, which utilized the Glauber-Sudarshan P-representation [11, 12] to simulate moments for classical fields, is extended to the positive P-representation, which can treat any quantum state. An extension to binning in multiple dimensions is also provided, as these are more sensitive tests of quantum advantage than one-dimensional distributions.
We show that in the challenging prospect of using quantum computers to demonstrate quantum computational advantage, the GBS approach has a very important advantage. Unlike many other proposed quantum computing architectures, there are multiple quantitative statistical tests that are dependent on the experimental parameters, which can be used to validate the output distributions of the data. If experimental data fail such tests, it is invalid. This is an important criterion, used in comparable tests of classical random number generators and cryptographic security [13, 14]. Implementing such tests requires algorithms that compute the correlations or moments without needing to compute the full #P hard distribution.
@@
An efficient and scalable method for validating GBS with threshold detectors, demonstrated for up to $16,000$ modes, was proposed in Ref.[15] and used to analyze data from a $100$-mode experiment [3]. The normally ordered positive P-representation was utilized to simulate grouped count probabilities (GCPs). This is a generalized binned photon counting distribution that includes all measurable marginal and binned correlations. The exponential sparsity of the detected count pattern probabilities means the full distribution cannot be either computed or measured. Therefore, validation tests based on marginals and binned probability distributions appear to be the most practical route to comprehensive validation.
@@
Like phase-space methods for threshold detectors [15], the PNR method is efficient and scalable. Increased dimension of binning causes an increase in computation time, where in the limit of dimension equaling mode number the GCP converges to the Hafnian. PNR phase-space methods allow one to perform simulations of either the ground truth distribution or modifications such as decoherence, measurement errors and shot-to-shot fluctuations. Such errors are typically not analyzed, despite evidence suggesting that a thermalized, non-ideal ground truth distribution provides stronger claims of quantum advantage than the ideal ground truth distribution [16]. Unlike threshold detection, PNR detector binning has the advantage of not requiring the computation of multidimensional inverse discrete Fourier transforms that increase computation time for multidimensional GCPs with threshold detection.
@@
The positive-P phase-space distribution is exact, non-singular and positive for any quantum state. The moments of the distribution are identical to normally ordered operator moments including intensity correlations [17], which are reproduced to any order. This property, coupled with probabilistic random sampling, makes the positive P-representation ideal for validating the statistics of nonclassical photon counting experiments such as GBS. It does not have the computational restraints imposed by other validation methods [18, 19, 20], since it is many orders of magnitude faster than direct classical simulation, with negligible rounding errors. While there are sampling errors, for current experimental parameters they are comparable to sampling errors in the measurement statistics. Both can be reduced in a controllable way.
@@
In order to determine the validity of using the positive P-representation to simulate nonclassical photon counting experiments with PNR detectors, we compare simulated GCP moments to the exactly known photon counting distributions for special cases. Here, multi-mode pure squeezed states are transformed by a Haar random unitary matrix which is multiplied by a uniform amplitude loss coefficient. For losses comparable to current experiments, we see excellent agreement. In the less realistic lossless case, the resulting non-Gaussian number distribution oscillates strongly between even and odd total counts. While this is far from current experiments, the standard positive P-representation requires a much larger number of phase-space samples to converge in such cases. However, this limit can also be efficiently simulated if necessary using a symmetry-projected [21] phase-space method, which will be treated elsewhere.
Finally, using photon count data from the recent Borealis experiment implementing PNR detectors [5], we apply the simulated GCPs to validate the moments of measured photon counting distributions by comparing theory and experiment. Differences are quantified using chi-square and Z-statistic tests. As with earlier work [15, 16], we show that the experimental GCPs are far from the ground truth distribution for pure squeezed state inputs. Instead, they are described much better by a different target distribution which includes decoherence and measurement error corrections of the transmission matrix. Although we focus specifically on implementations to GBS, we emphasize that one can use the positive P-representation to simulate GCPs of any linear photon counting experiment using classical or nonclassical photons.
This paper is structured as follows, Section II introduces the necessary background and theoretical methods for simulating GCPs of any photon counting experiment using PNR detectors with nonclassical photons in phase-space. Section IV then applies this method to the exactly known multi-mode squeezed state distributions, while Section V simulates the theoretical distributions corresponding to experimental data from the Borealis GBS claiming quantum advantage.
@@ II Grouped count probabilities
Since squeezed states are Gaussian, they are characterized entirely by their quadrature variances
$\left\langle\left(\Delta\hat{x}_{j}\right)^{2}\right\rangle$ $=2\left(n_{j}+m_{j}\right)+1=e^{2r_{j}}$
$\left\langle\left(\Delta\hat{y}_{j}\right)^{2}\right\rangle$ $=2\left(n_{j}-m_{j}\right)+1=e^{-2r_{j}},$ (2)
where the quadrature operators $\hat{x}_{j}=\hat{a}_{j}^{(\text{in})}+\hat{a}_{j}^{\dagger(\text{in})}$ and $\hat{y}_{j}=\left(\hat{a}_{j}^{(\text{in})}-\hat{a}_{j}^{\dagger(\text{in})}\right)/i$ satisfy the commutation relations $\left[\hat{x}_{j},\hat{y}_{k}\right]=2i\delta_{jk}$, and the mean photon number and coherence per mode is defined as $n_{j}=\left\langle\hat{a}_{j}^{\dagger(\text{in})}\hat{a}_{j}^{(\text{in})}\right\rangle=\sinh^{2}(r_{j})$ and $m_{j}=\left\langle\hat{a}_{j}^{(\text{in})}\hat{a}_{j}^{(\text{in})}\right\rangle=\cosh(r_{j})\sinh(r_{j})$, respectively.
While most theoretical analyses of GBS consider pure squeezed state inputs, this is impractical, since decoherence effects such as mode mismatches [3], partial distinguishability [23], and many more [24] will always be present during experimental applications in quantum optics. Therefore, a realistic model of input states are Gaussian thermalized squeezed states for which a simple parameterization was introduced in [15]. Here, a thermalization component $\epsilon$ alters the intensity of the input photons as measured by the coherence, which is reduced by a factor of $1-\epsilon$ such that $\tilde{m}_{j}=(1-\epsilon)m_{j}$. The quadrature variances are then altered as [15]
$\sigma_{x_{j}}^{2}=\left\langle\left(\Delta\hat{x}_{j}\right)^{2}\right\rangle$ $=2\left(n_{j}+\tilde{m}_{j}\right)$
$\sigma_{y_{j}}^{2}=\left\langle\left(\Delta\hat{y}_{j}\right)^{2}\right\rangle$ $=2\left(n_{j}-\tilde{m}_{j}\right),$ (3)
where for $\epsilon=0$ the pure squeezed state definition Eq.(2) returns, while $\epsilon=1$ produces fully decoherent thermal states with $\sigma_{x_{j}}^{2}=\sigma_{y_{j}}^{2}=2n_{j}$. More complex thermalized models are also possible, for example with mode dependent thermalization, $\epsilon_{j}$.
@@
$\hat{a}_{k}^{\dagger(\text{out})}=\sum_{j=1}^{N}T_{kj}\hat{a}_{j}^{\dagger(\text{in})}+\sum_{j=1}^{M}B_{kj}\hat{b}_{j}^{\dagger(\text{in})},$ (4)
where $\boldsymbol{B}$ is a random noise matrix representing photon loss to $j$-the environmental mode $\hat{b}_{j}^{\dagger(\text{in})}$. The resulting output photon number $\hat{n}^{\prime}_{k}=\hat{a}_{k}^{\dagger(\text{out})}\hat{a}_{k}^{(\text{out})}$ is then measured by a series of photodetectors.
@@ II.1 Photon counting projectors
where ::$\cdots$:: ⋯ : denotes normal ordering and $c_{j}=0,1,2,3,\dots$ is the number of measured photon counts at the $j$-th detector. Although theoretically there is no upper bound on $c_{j}$, practically PNR detectors saturate for some maximum number of counts $c_{j}^{(\text{max})}$. This value is variable, typically depending on the physical implementation of the detector.
@@
In the current generation of GBS networks, photon loss is the dominant noise source. Although there is evidence that sampling from a lossy GBS distribution is also a #P-hard computational task [31], output photons from experiments implementing high loss networks become classical [32], allowing the output distribution to become efficiently sampled. There is currently no clear boundary between the computationally hard and efficiently sampled regimes, but a classical algorithm that exploits large experimental photon loss has recently been developed [33] and shown to produce samples that are closer to the ideal ground truth than recent experiments claiming quantum advantage. The effects of thermalization on the computational complexity of GBS is currently unknown, despite thermalization causing nonclassical photons to become classical at a faster rate than pure losses.
@@ II.2 Grouped count probabilities for PNR detectors
Following from the nomenclature introduced in [15], we define the $d$-dimensional grouped photon counting distribution as the probability of observing $\boldsymbol{m}=[m_{1},\dots,m_{d}]$ grouped counts in $j$ distinct subsets of output modes $S_{j}=\{1,\dots,M/d\}$, such that
@@
Multi-dimensional probabilities are defined over a vector of disjoint subsets $\boldsymbol{S}=\text{$\left(S_{1},\dots,S_{d}\right)$}$ of mode numbers $\boldsymbol{M}=\left(M_{1},\dots,M_{d}\right)$, in which case GCPs become joint probability distributions. This is readily illustrated when $d=2$ and $n=M$, giving $\mathcal{G}_{S_{1},S_{2}}^{(M)}(m_{1},m_{2})$. Increasing the dimension to $d=M$ while keeping the correlation order constant at $n=M$ causes the GCP to converge to the Hafnian.
Therefore, if the correlation order satisfies $n=M$, an increase in GCP dimension corresponds to observing the $d$-th binned Glauber intensity correlation, hence high-order correlations become more statistically significant in the limit $d\rightarrow M$. If such correlations are present in the experimental data, multi-dimensional GCPs provide a more thorough test of quantum advantage as they allow one to directly compare these high-order correlations with their theoretical values [16]. If $n<M$, the GCP becomes a marginal probability as the remaining $M-n$ output modes are ignored.

### v2

@@ Abstract
An important challenge with the current generation of noisy, large-scale quantum computers is the question of validation. Does the hardware generate correct answers? If not, what are the errors? This issue is often combined with questions of computational advantage, but it is a fundamentally distinct issue. In current experiments, complete validation of the output statistics is generally not possible because it is exponentially hard to do so. Here, we apply phase-space simulation methods to partially verify recent experiments on Gaussian boson sampling (GBS) implementing photon-number resolving (PNR) detectors. The positive-P phase-space distribution is employed, as it uses probabilistic sampling to reduce complexity. It is $\sim 10^{18}$ times faster than direct classical simulation for experiments on $288$ modes where quantum computational advantage is claimed. When combined with binning and marginalization to improve statistics, multiple validation tests are efficiently computable, of which some tests can be carried out on experimental data. We show that the data as a whole shows discrepancies with theoretical predictions for perfect squeezing. However, a small modification of the GBS parameters greatly improves agreement for some tests. We suggest that such validation tests could form the basis of feedback methods to improve GBS experiments.
@@ I Introduction
Theoretical and experimental interest in photonic networks employed as quantum computers has grown considerably [1, 2]. These devices interfere photons from nonclassical states in a large scale interferometric network, then measure the resulting output photons using photodetectors. Of the proposed network architectures, Gaussian boson sampling (GBS) has proved scalable to the sizes needed to infer quantum advantage, which recent experiments have now claimed [3, 4, 5, 6]. In this case, photons are prepared in a squeezed state and the computational task is sampling from an output distribution with probabilities in a #P-hard computational class [7, 8]. For photon-number resolving (PNR) detectors this is a Hafnian, while for threshold detectors it is a Torontonian. Neither function is computable in polynomial time, so that the random counts cannot be classically replicated at large scale.
@@
A crucial question with all current generation quantum computers is how to test for errors in the computed outputs. An efficient and scalable method for validating GBS with threshold detectors, demonstrated for up to $16,000$ modes, was proposed in Ref.[9] and used to analyze data from a $100$-mode experiment [3]. The normally ordered positive-P phase-space representation was utilized to simulate grouped count probabilities (GCPs). This is a generalized binned photon counting distribution that includes all measurable marginal and binned correlations. The exponential sparsity of the detected count pattern probabilities means the full distribution cannot be either computed or measured. Therefore, validation tests based on marginals and binned probability distributions are the most practical method for validation, as is the case for other random number generators [10].
In this paper, we use positive P-representation simulations to test the validity of Gaussian boson samplers with photon number resolving (PNR) detectors, extending the tests that can be applied to experimental data. This is essential, as a number of large scale GBS experiments have claimed quantum advantage using either PNR [5] or threshold detectors [3, 4, 6]. To perform the tests, we generalize a binning method developed by Mandel [11, 12] for simulating photon statistics of classical light. The original formulation, which utilized the Glauber-Sudarshan P-representation [13, 14] to simulate moments for classical fields, is extended to the positive P-representation, which can treat any quantum state. An extension to binning in multiple dimensions is also provided, as these are more sensitive tests than one-dimensional distributions.
We show that in the challenging prospect of using quantum computers to demonstrate quantum computational advantage, the GBS approach has a very important advantage. Unlike many other proposed quantum computing architectures, there are multiple quantitative statistical tests that are dependent on the experimental parameters, which can be used to validate the output distributions of the data. Experimental data may fail such validation due to errors, potentially allowing classical simulations of comparable accuracy. This is an important criterion, similar to tests of classical random number generators and cryptographic security [15, 16]. Implementing a test suite requires algorithms that compute the correlations or moments without requiring the full #P hard distribution.
@@
Like phase-space methods for threshold detectors [9], the PNR method is efficient and scalable. Increased dimension of binning causes an increase in computation time, where in the limit of dimension equaling mode number the GCP converges to the Hafnian. PNR phase-space methods allow one to perform simulations of either the ground truth distribution or modifications of this distribution that include decoherence, measurement errors and shot-to-shot fluctuations. Such errors are typically not analyzed, despite evidence suggesting that a thermalized, non-ideal ground truth distribution provides stronger claims of quantum advantage than the ideal ground truth distribution [17]. Unlike threshold detection, PNR detector binning has the advantage of not requiring the computation of multidimensional inverse discrete Fourier transforms, which can increase computation time for increased dimension.
@@
The positive-P phase-space distribution is exact, non-singular and positive for any quantum state. The moments of the distribution are identical to normally ordered operator moments including intensity correlations [18], which are reproduced to any order. This property, coupled with probabilistic random sampling, makes the positive P-representation ideal for validating the statistics of nonclassical state photon counting experiments such as GBS. It does not have the computational restraints imposed by other validation methods [19, 20, 21], since it is many orders of magnitude faster than direct classical simulation, with negligible rounding errors. While there are sampling errors, for current experimental parameters they are comparable to sampling errors in the measurement statistics. Both can be reduced in a controllable way.
@@
In order to determine the validity of using the positive P-representation to simulate photon counting experiments with PNR detectors, we compare simulated GCP moments to the exactly known photon counting distributions for special cases. Here, multi-mode pure squeezed states are transformed by a Haar random unitary matrix which is multiplied by a uniform amplitude loss coefficient. For losses comparable to current experiments, we see excellent agreement. In the less realistic lossless case, the resulting non-Gaussian number distribution oscillates strongly between even and odd total counts. While this is far from current experiments, the standard positive P-representation requires a much larger number of phase-space samples to converge in such cases. However, this limit can also be efficiently simulated if necessary using a symmetry-projected [22] phase-space method, which will be treated elsewhere.
Finally, using photon count data from the recent Borealis experiment implementing PNR detectors [5], we apply the simulated GCPs to validate the moments of measured photon counting distributions by comparing theory and experiment. Differences are quantified using chi-square and Z-statistic tests. As with earlier work [9, 17], we show that the experimental GCPs are far from the ground truth distribution for pure squeezed state inputs. Instead, they are described much better by a different target distribution which includes decoherence and measurement error corrections of the transmission matrix. Although we focus specifically on implementations to GBS, we emphasize that one can use the positive P-representation to simulate GCPs of any linear photon counting experiment using classical or nonclassical state photons.
This paper is structured as follows, Sections II and III introduce the necessary background and theoretical methods for simulating GCPs of any photon counting experiment using PNR detectors in phase-space. Section IV then applies this method to the exactly known multi-mode squeezed state distributions, while Section V simulates the theoretical distributions corresponding to experimental data from the Borealis GBS claiming quantum advantage. Finally, VI summarizes the main findings of this paper.
@@ II Grouped count probabilities
Since squeezed states are Gaussian, they are characterized by their quadrature variances
$\left\langle:\left(\Delta\hat{x}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}+m_{j}\right)=e^{2r_{j}}-1$
$\left\langle:\left(\Delta\hat{y}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}-m_{j}\right)=e^{-2r_{j}}-1,$ (2)
where the quadrature operators $\hat{x}_{j}=\hat{a}_{j}^{(\text{in})}+\hat{a}_{j}^{\dagger(\text{in})}$ and $\hat{y}_{j}=\left(\hat{a}_{j}^{(\text{in})}-\hat{a}_{j}^{\dagger(\text{in})}\right)/i$ satisfy the commutation relations $\left[\hat{x}_{j},\hat{y}_{k}\right]=2i\delta_{jk}$, and the mean photon number and coherence per mode is defined as $n_{j}=\left\langle\hat{a}_{j}^{\dagger(\text{in})}\hat{a}_{j}^{(\text{in})}\right\rangle=\sinh^{2}(r_{j})$ and $m_{j}=\left\langle\hat{a}_{j}^{(\text{in})}\hat{a}_{j}^{(\text{in})}\right\rangle=\cosh(r_{j})\sinh(r_{j})$, respectively. Throughout this paper, we use ::$\cdots$:: ⋯ : to denote normal ordering.
While most theoretical analyses of GBS consider pure squeezed state inputs, this is impractical, since decoherence effects such as mode mismatches [3], partial distinguishability [24], and many more [25] will always be present during experimental applications in quantum optics. Therefore, a realistic model of input states are Gaussian thermalized squeezed states for which a simple parameterization was introduced in [9]. Here, a thermalization component $\epsilon$ alters the intensity of the input photons as measured by the coherence, which is reduced by a factor of $1-\epsilon$ such that $\tilde{m}_{j}=(1-\epsilon)m_{j}$. The quadrature variances are altered as [9]
$\sigma_{x_{j}}^{2}=\left\langle:\left(\Delta\hat{x}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}+\tilde{m}_{j}\right)$
$\sigma_{y_{j}}^{2}=\left\langle:\left(\Delta\hat{y}_{j}\right)^{2}:\right\rangle$ $=2\left(n_{j}-\tilde{m}_{j}\right),$ (3)
where for $\epsilon=0$, the pure squeezed state definition Eq.(2) returns, while $\epsilon=1$ produces fully decoherent thermal states with $\sigma_{x_{j}}^{2}=\sigma_{y_{j}}^{2}=2n_{j}$. More complex thermalized models are also possible, for example with mode dependent thermalization, $\epsilon_{j}$.
@@
$\hat{a}_{k}^{\dagger(\text{out})}=\sum_{j=1}^{N}T_{kj}\hat{a}_{j}^{\dagger(\text{in})}+\sum_{j=1}^{M}L_{kj}\hat{b}_{j}^{\dagger(\text{in})},$ (4)
where $\boldsymbol{L}$ is a random noise matrix representing photon loss to $j$-the environmental mode $\hat{b}_{j}^{\dagger(\text{in})}$. The resulting output photon number $\hat{n}^{\prime}_{k}=\hat{a}_{k}^{\dagger(\text{out})}\hat{a}_{k}^{(\text{out})}$ is then measured by a series of photodetectors.
@@ II.1 Photon counting projectors
where $c_{j}=0,1,2,3,\dots$ is the number of measured photon counts at the $j$-th detector. Although theoretically there is no upper bound on $c_{j}$, practically PNR detectors saturate for some maximum number of counts $c_{j}^{(\text{max})}$. This value is variable, typically depending on the physical implementation of the detector.
@@
In the current generation of GBS networks, photon loss is the dominant noise source. Although there is evidence that sampling from a lossy GBS distribution is also a #P-hard computational task [32], output photons from experiments implementing high loss networks become classical [33], allowing the output distribution to become efficiently sampled. There is currently no clear boundary between the computationally hard and efficiently sampled regimes, but a classical algorithm that exploits large experimental photon loss has recently been developed [34] and shown to produce samples that are closer to the ideal ground truth than recent experiments claiming quantum advantage. The effects of thermalization on the computational complexity of GBS is currently unknown, despite thermalization causing photons from nonclassical states to become classical at a faster rate than pure losses.
@@ II.2 Grouped count probabilities for PNR detectors
Following from the nomenclature introduced in [9], we define the $d$-dimensional grouped photon counting distribution as the probability of observing $\boldsymbol{m}=[m_{1},\dots,m_{d}]$ grouped counts in $j=1,\dots,d$ distinct subsets of output modes $\boldsymbol{S}=\text{$\left(S_{1},\dots,S_{d}\right)$}$, such that
@@
The correlation order and dimension determines the type of observable one is simulating. If $n=M$, all correlation orders are simulated in phase-space. Meanwhile $n<M$ denotes a marginal probability, since the remaining $M-n$ output modes are ignored. If $n=M$ and $d=1$, the GCP becomes the univariate distribution $\mathcal{G}_{S}^{(M)}(m)$ where all modes are contained within a single subset $S=\{1,\dots,M\}$. This is called the total count distribution as it’s the probability of observing $m$ counts in any pattern and is identical to the total photon-number distribution of Ref. [5].
When $d>1$, multi-dimensional GCPs are simulated. These are joint probability distributions over $\boldsymbol{m}$ counts. For example, one of the $d=2$ GCPs is the bivariate photon count distribution $\mathcal{G}_{S_{1},S_{2}}^{(M)}(m_{1},m_{2})$, which is the probability of observing $m_{1}$ counts in $S_{1}=\{1,\dots,M_{1}\}$, and $m_{2}$ counts in $S_{2}=\{M_{1}+1,\dots,M\}$. The largest possible dimension is $d=M$ where the vector of subsets becomes $\boldsymbol{S}=(S_{1},S_{2},\dots,S_{M})=(\{1\},\{2\},\dots,\{M\})$. In this case, the GCP to converges to the Hafnian as the joint probabilities have become pattern probabilities.
Provided one has the computational resources (see Section IV.3), comparisons of multi-dimensional GCPs provide a number of benefits over their $d=1$ counterparts. The first is the number of count bins scales as $\mathcal{B}=(Md^{-1}+1)^{d}$, allowing more detailed comparisons with theory and experiment. The second is by randomly permuting the measured count patterns, an exponentially large number of comparison tests can be performed on experimental data, where each permutation allows one to compare different binned correlation moments [17].
So long as $n=M$ all correlation orders are simulated in phase-space. However, since GCP probabilities converge to the Hafnian as $d\rightarrow M$, a greater range of distinct correlations are statistically significant in this limit. If these correlations are nonclassical, multi-dimensional GCPs provide a more thorough test of quantum advantage, because they allow one to directly compare more sums of high-order correlations with their theoretical values [17], and these tests are more sensitive to the exact squeezing and transfer matrix parameter values.